OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3. y - 2 = -3(x + 7) y + 7 = -3(x - 2) y + 2 = -3(x - 7) y - 7 = -3(x + 2)

OpenStudy (anonymous):

point slope form is y-y1=m(x-x1) so therefore your equation would be y-(-2)=-3(x-7) y+2=-3(x-7)

hartnn (hartnn):

the equation of line with slope m and point (x1,y1) is y-y1=m(x-x1)

OpenStudy (anonymous):

so the third choice.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

no problem:)